Find the value of the expression in terms of tan  : ( sin2x + sin4x )/ ( 1+cos2x + cos4x )

Grade:11

1 Answers

Lab Bhattacharjee
8 years ago
\sin2x+\sin4x=\sin2x(1+2\cos2x)\\ 1+\cos2x+\cos4x=\cos2x+2\cos^22x=\cos2x(1+2\cos2x)